OpenStudy (anonymous):

(-3X) + 40 = 38 3/2 2/3 -2/3 -3/2

OpenStudy (anonymous):

(-3x)+40=38 (-3x)=38-40 -3x=-2 solve x..

OpenStudy (anonymous):

2/3